OpenStudy (anonymous):

If two different integers divided by same divisor leaves remainder 7 and 9 respectively,and when their sum divided by same divisor it leaves remainder 5, then what is the divisor?

imqwerty (imqwerty):

ok let say that divisor=d nd the 2 integers are m and n so m=da+7 equation 1 n=db+9 equation 2 where a and b are quotients :) nd then u divide m+n u get- m+n=dq+5 and adding equation 1 and 2 u get- m+n=da+db+7+9 m+n=d(a+b)+16 but the remainder is 5 to get remainder = 5 d shuld be 11 :)